# Learning the Covariance of Treatment Effects Across Many Weak Experiments

### v1

@@ 5.1 LIMLK as Total Least Squares
We now consider the LIMLK estimator as an alternative to the naive estimator based on the covariance matrix of estimated treatment effects. As described in Section 4.3, LIMLK is equivalent to performing Total Least Squares (TLS) after applying the linear transformation $\Omega^{-1/2}$ to the scatterplot of estimated treatment effects. Intuitively, this is because applying this transformation on both sides to $\widehat{\Sigma}_{K}$ yields, in expectation, the transformed true covariance matrix $\Omega^{-1/2}\Lambda_{K}\Omega^{-1/2}$ plus isotropic noise – a classic error-in-variables setup. TLS is an effective method for addressing error-in-variables as isotropic noise does not change the eigenvectors of $\widehat{\Sigma}_{K}$.
Figure 3: Comparison of Naive, LIML, and Total Covariance Estimators Without Direct Effects
@@
5.2 Comparing the Naive, LIMLK, and TC Estimators
@@
To illustrate the effects of unit-level noise on our estimators, we simulate two surrogates $S_{1}$ and $S_{2}$ with corresponding structural parameters $\beta_{1}$ and $\beta_{2}$. $S_{1}$ is highly correlated with the primary outcome $Y$ in terms of the noise covariance matrix $\Omega$, while $S_{2}$ is uncorrelated with $Y$. As a result, we expect the naive estimator to perform relatively badly for $S_{1}$ (as it ignores the unit-level covariance) compared to $S_{2}$ and the LIMLK and TC estimators. To assess the robustness of our estimators to direct effects, we simulate data according to the linear structural models in Subsections 3.1 (no direct effects) and 3.2 (direct effects under INSIDE). The specific simulation parameters are described in the Appendix.
@@
As Figure 3 shows, the naive estimator is heavily biased. LIMLK and TC are substantially much more accurate than the naive estimator. Again, the relative accuracy of LIMLK and TC is more apparent for $\beta_{1}$, which has the greater distortion from the unit-level covariance. Without direct effects, LIMLK is more efficient compared to TC, with a considerably smaller standard error at smaller values of $n$.
@@
However, LIMLK is highly sensitive to the assumption of no direct effects. As Figure 4 shows, when we introduce direct effects, LIMLK is clearly inconsistent for $\beta$. In contrast, TC is robust and remains consistent for $\beta$ even when we add direct effects. However, note that, regardless of estimator, the interpretation of $\beta$ is sensitive to the presence of direct effects. If $S$ fully mediates the effect of $A$ on $Y$, $h(S)=\beta S$ is a surrogate index, and the ATE of a new treatment on $Y$ is equal to the ATE on $h(S)$. If $S$ does not fully mediate the effect of $A$ on $Y$, and direct effects follow the INSIDE assumption, then $\beta$ can be interpreted as the portion of the treatment effect of $Y$ that is mediated by $S$.
Figure 4: Comparison of Naive, LIML, and Total Covariance Estimators With Direct Effects
6 Linear Models of Treatment Effect Covariances at Netflix.
@@
This section briefly discusses the application of our methods to construct linear proxy metric indices for digital experiments at Netflix. Netflix runs thousands of experiments per year on many diverse parts of the business. Each area of experimentation is staffed by independent data science and engineering teams. While every team ultimately aims to lift the same north star metrics (e.g., long-term revenue), each has also developed secondary metrics that are more sensitive and practical to measure in short-term A/B tests (e.g., user engagement or latency). To complicate matters further, teams are constantly innovating on these secondary metrics to find the right balance of sensitivity and impact on the north star metrics.
@@
In this setting, linear models of relationships between treatment effects are a highly useful tool for coordinating these efforts and aligning them towards the same proxies for the north star:
  1. 1.
@@
  2. 2.
    Informing metrics innovation. To minimize wasted effort on metric development, it’s also important to understand how metrics correlate with the north star “net of” existing metrics. The linear models presented in this paper allow teams to quickly understand if an additional secondary metric is significantly correlated with north star metrics after partialling out their correlations with the secondary metrics that are already included in the proxy metric.
  3. 3.
@@
7 Conclusion
In this paper, we discuss the construction of proxy metrics via meta-analysis of many experiments. A useful parameter in this setting is the covariance matrix of treatment effects across metrics in the population of experiments. While estimating this parameter is computationally convenient, it presents challenges, especially when treatment effects are small relative to unit-level covariances. We present two estimators for overcoming these challenges and show how their estimands relate to structural parameters under different causal models.
@@ 7 Conclusion
Beyond the construction of proxy metrics, there are other opportunities to leverage short-term observations for inference on long-term outcomes. For example, Athey et al. (2020); Imbens et al. (2022); Chen and Ritzwoller (2023) consider how short-term observations from one randomized experiment can remove confounding in an observational study with long-term observations, and Kallus and Mao (2020) consider efficiency gains from including units with no long-term observations. Considering the use of many, albeit weak, experiments as instruments in these settings may uncover new opportunities to target long-term outcomes under weaker assumptions.

### v2

@@ 5.1. LIMLK as Total Least Squares
We now consider the LIMLK estimator as an alternative to the naive estimator based on the covariance matrix of estimated treatment effects. As described in Section 4.3, LIMLK is equivalent to performing Total Least Squares (TLS) after applying the linear transformation $\Omega^{-1/2}$ to the scatterplot of estimated treatment effects. Intuitively, this is because applying this transformation on both sides to $\widehat{\Sigma}_{K}$ yields, in expectation, the transformed true covariance matrix $\Omega^{-1/2}\Lambda_{K}\Omega^{-1/2}$ plus isotropic noise — a classic error-in-variables setup. TLS is an effective method for addressing error-in-variables as isotropic noise does not change the eigenvectors of $\widehat{\Sigma}_{K}$.
@@
5.2. Comparing the Naive, LIMLK, and TC Estimators
@@
To illustrate the effects of unit-level noise on our estimators, we simulate two surrogates $S_{1}$ and $S_{2}$ with corresponding structural parameters $\beta_{1}$ and $\beta_{2}$. $S_{1}$ is highly correlated with the primary outcome $Y$ in terms of the noise covariance matrix $\Omega$, while $S_{2}$ is uncorrelated with $Y$. As a result, we expect the naive estimator to perform relatively badly for $S_{1}$ (as it ignores the unit-level covariance) compared to $S_{2}$ and the LIMLK and TC estimators. To assess the robustness of our estimators to direct effects, we simulate data according to the linear structural models in Subsections 3.1 (no direct effects) and 3.2 (direct effects under INSIDE). [^3] Code to replicate our simulations and all figures except those in Section 6 can be found at https://github.com/winston-chou/linear-proxy-metrics.
@@
As Figure 3 shows, the naive estimator is heavily biased. LIMLK and TC are substantially much more accurate than the naive estimator. The relative accuracy of LIMLK and TC is more apparent for $\beta_{1}$, which has the greater distortion from the unit-level covariance. Without direct effects, LIMLK is more efficient compared to TC, with a considerably smaller standard error at smaller values of $n$.
@@
However, LIMLK is highly sensitive to the assumption of no direct effects. As Figure 4 shows, when we introduce direct effects, LIMLK is clearly inconsistent for $\beta$. In contrast, TC is robust and remains consistent for $\beta$. However, note that, regardless of estimator, the interpretation of $\beta$ is sensitive to the presence of direct effects. If $S$ fully mediates the effect of $A$ on $Y$, $h(S)=\beta S$ is a surrogate index, and the ATE of a new treatment on $Y$ is equal to the ATE on $h(S)$. If $S$ does not fully mediate the effect of $A$ on $Y$, and direct effects follow the INSIDE assumption, then $\beta$ can be interpreted as the portion of the treatment effect of $Y$ that is mediated by $S$.
Figure 3. Comparison of Naive, LIML, and Total Covariance Estimators Without Direct Effects. When there are no direct effects, such that the secondary outcomes fully mediate the effect of the intervention on the primary outcome, our LIMLK (limlk) and Total Covariance (tc) estimators are substantially less biased than the naive regression of the estimated ATEs on the primary outcome on the estimated ATEs on the secondary outcomes (naive). LIMLK has lower variance than TC, which is most apparent at small sample sizes.
Figure 4. Comparison of Naive, LIML, and Total Covariance Estimators With Direct Effects. When the intervention directly affects the primary outcome, LIMLK can be extremely biased. In contrast, TC is robust to direct effects, although the causal interpretation of its estimand changes if there are direct effects. The naive estimator is biased regardless of whether there are direct effects or not.
6. Linear Models of Treatment Effect Covariances at Netflix
Figure 5. How Our Method Improves Treatment Effect Covariance Estimation at Netflix. This figure shows how correlated measurement error exaggerates the correlation between estimated treatment effects, especially at smaller sample sizes (top), and how our TC estimator reduces this bias by subtracting a scaled estimate of the measurement error (bottom).
This section discusses the real-world application of our methods to construct linear proxy metric indices for experimentation at Netflix. Netflix has a sophisticated experimentation platform that runs thousands of experiments on millions of experimental units (forsell2020success, ). Still, because the signal-to-noise ratio of each experiment is small, measurement error poses challenges. This is shown in the top panel of Figure 5, which plots the correlation between estimated treatment effects on one Short-term Metric and one Long-term Metric across 96 arbitrarily-chosen treatment-control comparisons. The Short- and Long-term Metrics are highly correlated across experimental units, introducing correlated measurement error that exaggerates the correlation across treatment effects. To demonstrate the effect of experiment size on this bias, we show the scatterplot of estimated treatment effects after downsampling each experiment to one million units (white circles) and 15 million units (gray circles). The slope of the OLS in the former scatterplot is about 50% larger than the equivalent slope in the latter scatterplot. Note that we have hidden the x- and y-axis tick marks in this figure for confidentiality reasons.
In the bottom panel of Figure 5, we plot the absolute bias of the observed covariance between the Short- and Long-term Metrics at various subsample sizes, using as ground truth the TC-corrected covariance across non-downsampled experiments. We also plot the absolute bias of the TC-corrected covariance for comparison. The median absolute bias reduction from applying our TC estimator is substantial — approximately 63%.
@@ 6. Linear Models of Treatment Effect Covariances at Netflix
In addition to these empirical benefits, our method is also well-suited to the decentralized and rapidly-evolving practice of experimentation at Netflix. Netflix runs thousands of experiments per year on many diverse parts of the business. Each area of experimentation is staffed by independent data science and engineering teams. While every team ultimately aims to lift the same north star metrics (e.g., long-term revenue), each has also developed secondary metrics that are more sensitive and practical to measure in short-term A/B tests (e.g., user engagement or latency). To complicate matters further, teams are constantly innovating on these secondary metrics to find the right balance of sensitivity and impact on the north star metrics.
@@
In this decentralized environment, linear models of relationships between treatment effects are a highly useful tool for coordinating these efforts and aligning them towards the same proxies for the north star:
  1. (1)
@@
  2. (2)
    Informing metrics innovation. To minimize wasted effort on metric development, it is also important to understand how metrics correlate with the north star “net of” existing metrics.
  3. (3)
@@
7. Conclusion
In this paper, we discuss the construction of proxy metrics via meta-analysis of many experiments. A useful parameter in this setting is the covariance matrix of treatment effects across metrics in the population of experiments. While estimating this parameter is computationally convenient, it presents challenges, especially when treatment effects are small relative to unit-level noise. We present estimators for overcoming these challenges and show how their estimands relate to structural parameters under different causal models.
@@ 7. Conclusion
Beyond the construction of proxy metrics, there are other opportunities to leverage short-term observations for inference on long-term outcomes. For example, (athey2020combining, ; imbens2022long, ; chen2023semiparametric, ) consider how short-term observations from one randomized experiment can remove confounding in an observational study with long-term observations, and (kallus2020role, ) consider efficiency gains from including units with no long-term observations. Considering the use of many, albeit weak, experiments as instruments in these settings may uncover new opportunities to target long-term outcomes under weaker assumptions.